OpenStudy (anonymous):

I will give you a medal and be your fan! What is the difference? 1/6 - 5/8 A. -19/24 B. -11/24 C. 11/24 D. 2

OpenStudy (anonymous):

B

OpenStudy (anonymous):

-11/24

OpenStudy (anonymous):

Can you explain how you got the answer?

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

you subtract the numbers then simplify

OpenStudy (anonymous):

L.C.D of 6 and 8 is 24 \[\frac{ 1 }{ 6 }-\frac{ 5 }{ 8 }=\frac{ 4-15 }{ 24 }=?\]

OpenStudy (anonymous):

-11/24

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Thank you guys for the help

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Yup dont worry about it